Bobby has a bond earning interest 11% yearly and another worth 3 times as much earning 10% simple inteest. If the total yearly interest is $307.50, find the value of each bond.
Answer by ptaylor(2198)

Let x=value of 11% bond
Then 3x=value of 10% bond
Interest(I)=Principal(P) times Rate(R) times Time(T) or I=PRT; T=1
So our equation to solve is:
0.11x+0.10(3x)=$307.50
0.11x+0.3x=307.50
0.41x=307.50 divide both sides by 0.41
x=$750-------------------------------------value of 11% bond
3x=3(750)=$2250-------------------------------value of 10% bond
CK
750(.11)+2250(.10)=307.50
82.5+225=307.50
307.50=307.50

Hope this helps----ptaylor